OpenStudy (anonymous):

Find the surface area of a sphere with a volume of 36π in^3.

OpenStudy (goldphenoix):

To find the surface area of a sphere, you want to use the formula: 4π r^2 We are given with the volume of a sphere. The volume of the sphere is: 36π We need to find the radius of the sphere to get the surface area, correct?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Oh yeah, to find the volume of a sphere, you use the formula: (4/3) π r^3 So: (4/3) π r^3 = 30π We want to find r. So what do we do first?

OpenStudy (anonymous):

cancel out pi on both sides

OpenStudy (goldphenoix):

Alright. So we would get: (4/3) r^3 = 30 Now what would we do?

OpenStudy (anonymous):

divide both sides by 4/3?

OpenStudy (goldphenoix):

Hmm. We want to get rid of (4/3). So (4/3) divided by (4/3) cancel out?

OpenStudy (anonymous):

yup

OpenStudy (goldphenoix):

Okay. So what does 30 divided by (4/3) give you?

OpenStudy (anonymous):

22.5

OpenStudy (goldphenoix):

Hmm. Alright. So we have: r^3 = 22.5 Now what would we do?

OpenStudy (anonymous):

this is the part i dont get. if it was r^2 you would just find the square root of 22.5 but i dont know how to do it with r^3

OpenStudy (goldphenoix):

Well. If you want to get rid of square, then you would square root it. If you want to get rid of cube, then you would have to cube root it.

OpenStudy (anonymous):

so how do you do that

OpenStudy (goldphenoix):

It would look like: \[\large \large \large \sqrt[3]{22.5}\] You can use a calculator if you want.

OpenStudy (anonymous):

so the answer is about 2.82

OpenStudy (goldphenoix):

Yes. Correct. So the radius is 2.82, approximately. Now to find the surface area of the sphere, you use the formula: 4π r^2 All you need to do is plug it in. After plugging it in, your equation should look like: 4π * 2.82^2

OpenStudy (anonymous):

oh okay that makes sense! thank you!

OpenStudy (goldphenoix):

No problem. :) I hope this help!